Keeping the total length of coil unchanged, the no. Of rounds in the coil are doubled. Its new self inductance will be 1.  Four times. 2.  Double 3. Half. 4. Unchanged 

Grade:12

1 Answers

ASHISH KUMAR
2 years ago
magnetic field for solenoid = B = \muNI 
and formula for self inductance is
     N\phi = L I  ( where L is self inductance )
and \phi(flux) = BACos \theta  ( \theta = 0 so Cos \theta= 1)
 case 1 )  N ( \muNI * A) = LI 
               L = \mu N^2 A
case 2)  now no of turn is doubled
so 2N ( \mu(2N) I *A ) = L' I 
  L' = 4 \muN^2 A
hence  Its new self inductance will be four times
option 1 Four times is correct